OpenStudy (anonymous):

how to solve this indeterminate form ( use L'hospital) , please show me the steps

OpenStudy (anonymous):

\[\lim_{x \rightarrow o}( \sin x \log x^2)\]

ganeshie8 (ganeshie8):

\[\large \lim_{x \rightarrow 0} \left( \sin x \log x^2\right) \] \[\large \lim_{x \rightarrow 0} \left( \dfrac{\sin x}{x} \right)\times \lim_{x \rightarrow 0}\left(x\log x^2\right) \]

ganeshie8 (ganeshie8):

\[\large \lim_{x \rightarrow 0}\left(x\log x^2\right) \]

ganeshie8 (ganeshie8):

\[\large \lim_{x \rightarrow 0}\left( \dfrac{\log x^2}{1/x}\right) \]

ganeshie8 (ganeshie8):

use L'Hopital now

ganeshie8 (ganeshie8):

\[\large \lim_{x \rightarrow 0}\left( \dfrac{1/x^2 (2x)}{-1/x^2}\right) \] \[\large \lim_{x \rightarrow 0}\left(-2x\right) \] \[\large 0\]

ganeshie8 (ganeshie8):

see if that makes more or less sense..

OpenStudy (anonymous):

you are amazing @ganeshie8 .. made the sum look so easy :) ..thanks a lot !!

ganeshie8 (ganeshie8):

yw :)